A particle is projected from the ground in a vertical direction at t=0. At t=0.8sec, it reaches h=14m. It will again come to the same height at t=? [g=10m/s2

Grade:11

1 Answers

Arun
6 years ago
Dear Ayush
 
h= v₀•t-g•t²/2 , 
v₀=h/t +gt/2 = 14/0.8 +9.8•0.8/2 = 21.42 m/s. 
Time of the motion to the highest point 
t₀=v₀/g =21.42/9.8=2.19 s. 
t1=2•2.19 – 0.8=3.58 s.